OpenStudy (anonymous):

The students in a small class are divided into three teams A, B and C. Each week two of the teams are selected at random to participate in a competition. What is the probability that team C is selected at least two times during the next three weeks?

OpenStudy (kropot72):

The probability that team C is selected in a random draw is given by \[P(team\ C\ selected)=\frac{\left(\begin{matrix}1 \\ 1\end{matrix}\right)\left(\begin{matrix}2 \\ 1\end{matrix}\right)}{\left(\begin{matrix}3 \\ 2\end{matrix}\right)}=\frac{2}{3}\] Now we have the probability of team C being selected, the probability that team C is selected at least two times during the next three weeks can be found from the binomial distribution. Do you know how to continue from here?

OpenStudy (anonymous):

Not really :( I read the answer similar to this one that you posted but I still don't get it!

OpenStudy (kropot72):

We can use the binomial distribution to find the probability of team C being selected 3 times during the next 3 weeks as follows: \[P(selected\ 3\ out\ of\ 3)=\left(\begin{matrix}3 \\ 3\end{matrix}\right)\times (\frac{2}{3})^{3}\times (\frac{1}{3})^{0}=(\frac{2}{3})^{3}=you\ can\ calculate\] When you have calculated that probability, subtract it from 1 to find the required probability.

OpenStudy (kropot72):

@HelpMeWithMath5 Are you there?

OpenStudy (anonymous):

Yes thank you! the answer I got was 0.7037, but when I entered that answer into the online program for one of my assignments, it says its wrong :S

OpenStudy (kropot72):

How many decimal places were required for the answer?

OpenStudy (anonymous):

4 decimal places

OpenStudy (kropot72):

Sorry, my bad. I did not read the question carefully enough. To get the correct answer the probabilities of two selections and 3 selections must be added. \[P(selected\ 2\ out\ of\ 3)=\left(\begin{matrix}3 \\ 2\end{matrix}\right) \times (\frac{2}{3})^{2}\times (\frac{1}{3})=3\times (\frac{2}{3})^{2}\times (\frac{1}{3})=you\ can\ calculate\]

OpenStudy (kropot72):

@HelpMeWithMath5 How are you doing?

OpenStudy (anonymous):

Wow something happened to my computer and it automatically sent in nothing as an answer and that was my final try, A message popped up and said the answer was 0.7407!!! This is very strange

OpenStudy (anonymous):

But thank you anyway!!!

OpenStudy (kropot72):

The correct answer is indeed 0.7407 P(2 out of 3) + P(3 out of 3) = 0.444444444 + 0.296296296

OpenStudy (anonymous):

Yes I got that, I just emailed my teacher the answer and she is gonna fix my score. Thanks for the help!

OpenStudy (kropot72):

You're welcome :)